OpenStudy (anonymous):

Adam has $5.25 in nickels and dimes. If he has 15 more dimes than nickels, how many nickels does he have?

OpenStudy (anonymous):

Option's are..... 27 33 30 25 20

OpenStudy (anonymous):

25 nickels

OpenStudy (anonymous):

and 40 dimes :)

OpenStudy (anonymous):

.10d + .05n = 5.25 d = n + 15 now sub in n + 15 in for d in the 1st equation .10(n + 15) + .05n = 5.25 -- distribute through the parenthesis .10n + 1.50 + .05n = 5.25 -- subtract 1.50 from both sides .10n + .05n = 5.25 - 1.50 .15n = 3.75 n = 3.75/.15 n = 25 =====> 25 nickels now sub 25 in for n in either of the equations to find d d = n + 15 d = 25 + 15 d = 40 ====> 40 dimes I know you already were given the answer....I just wanted to show you how to do it :)

OpenStudy (anonymous):

.10d + .05n = 5.25 how do get this....@kelliegirl33

OpenStudy (anonymous):

Now you know exactly how to do it. :) I should have done that

OpenStudy (anonymous):

that's the formula you use. Or at least how you set it up

OpenStudy (anonymous):

dimes are worth .10 and nickels are worth .05....the variable " n " stands for number of coins. So....10d + .05n = 5.25

OpenStudy (anonymous):

because the currency is times by its value to obtain its real value

OpenStudy (anonymous):

just listen to them ;)

OpenStudy (anonymous):

for example, u have x n-value coins, in order to get the value of the coins, u take x multiply by n, by logic

OpenStudy (anonymous):

you understand chetan ?

OpenStudy (anonymous):

kellie, will u help me with a qn? i have loads trouble doin it. pls, i need to finish it by tmr

OpenStudy (anonymous):

if you were given quarters instead of dimes, it would be .25n because quarters are worth 25 cents

OpenStudy (anonymous):

but thax for your explnation....@kelliegirl33 @trollwillcontinue @VIbarguen1

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

u are always welcome =)

OpenStudy (anonymous):

:)